OpenStudy (anonymous):

find the slope of the line through the points (9,-4) and (9,5) 1 0 undefined 9

hero (hero):

\[m = \frac{y_2 - y_1}{x_2 - x_1}\]

hero (hero):

That's the slope formula by the way where m = slope \[(9,4) = (x_1, y_1) \space\text {and} (9,5) = (x_2,y_2)\]

myininaya (myininaya):

And it really doesn't matter if it is (9,4) or (9,-4)

hero (hero):

oops

hero (hero):

where's the edit button?

myininaya (myininaya):

Great work hero.

OpenStudy (anonymous):

i'm sorry, so what is it?

myininaya (myininaya):

Use the slope formula Hero gave you and plug in \[x_1=9,y_1=-4,x_2=9, y_2=5\]

myininaya (myininaya):

In other words replace x1 with 9 replace y1 with -4 replace x2 with 9 replace y2 with 5 and then simplify

myininaya (myininaya):

But what happens in the denominator ?

OpenStudy (anonymous):

1?

myininaya (myininaya):

hmmm so you have 9-9 in the denominator since both x's are 9

myininaya (myininaya):

9-9=0 right?

myininaya (myininaya):

But when we have division by zero we call that what?

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

slope m = (5 -4)/ ( 9 -9) = 1/0 -> slope is undefined ( vertical line x = c)